Hi, I have to determine the linear function f(t) with slope -1 and f(2)= -1.
I have done slope -1 and (2, -1).
form f(x) =ax+b
a= -1
f(x) = -1
Since f(2)=1
f(2)= -1(2)+b
2=-2+b
2+2=b
4=b
f(t) = -1t+4
Is this even remotely on the right track??
thanks for your time!

Answer by rapaljer(4671)

You are almost right. You made a couple of very small errors, which I have corrected below for you.

I have done slope -1 and (2, -1).
form f(x) =ax+b
a= -1
f(x) = -1
Since f(2)=-1
f(2)= -1(2)+b
-1=-2+b
-1+2=b
1=b
f(t) = -1t+1

Now, check it! The slope is indeed -1.
f(2)= -1(2) + 1 = -2 +1 = -1, as it should be.

R^2 at SCC